OpenStudy (anonymous):

f open x close equals x squared minus 5 x minus 6 ,

OpenStudy (ahsome):

Equation: \[f(x)=x^2-5x-6\]

OpenStudy (anonymous):

what is one of the zeros?

OpenStudy (ahsome):

One of the zero?

OpenStudy (anonymous):

Given the function, f open x close equals x squared minus 5 x minus 6 , what is one of the zeros? A: x equals 6 B: x equals 3 C: x equals 1 D: x equals minus 6

OpenStudy (ahsome):

If the answer is right, then when you sub the value into the equation is 0 Example is:\[x^2−5x−6\] Sub in \(x=6\) \[6^2-5*6-6\] \[36-30-6=0\] Therefore, \(x=6\) is a factor of the the equation.

OpenStudy (anonymous):

thank you!

OpenStudy (ahsome):

No problem @dcedge99 :)